Consider the following numerical example of simple Keynesian model with no government sending , taxes or foreign sector.
C=100+0,9y
I=50
A. What is the value of the marginal prosperity to consume, in this model?
B. Use a graph to illustrate the equilibrium level of output?
C.Calculate the equilibrium level of output?
D. In equilibrium what is the value of consumption spending?use this number to verify that the sum of C and I in equilibrium equals the Value for equilibrium output you obtained above
1
Expert's answer
2019-08-14T09:25:31-0400

A. there is no change in consumption so MPC=0


B.



C.C=100+0,9y, I=50 , Y=AD, Y=C+I

Y=100+0.9Y+50, Y=1500


D. C=100+0.9(1500)= 1450

The sum of C and I inthe equlibrium = 1450+50=1500